### v1

# CustomListener: Text-guided Responsive Interaction for User-friendly Listening Head Generation

Abstract
Listening head generation aims to synthesize a non-verbal responsive listener head by modeling the correlation between the speaker and the listener in dynamic conversion. The applications of listener agent generation in virtual interaction have promoted many works achieving the diverse and fine-grained motion generation. However, they can only manipulate motions through simple emotional labels, but cannot freely control the listener’s motions. Since listener agents should have human-like attributes (e.g. identity, personality) which can be freely customized by users, this limits their realism. In this paper, we propose a user-friendly framework called CustomListener to realize the free-form text prior guided listener generation. To achieve speaker-listener coordination, we design a Static to Dynamic Portrait module (SDP), which interacts with speaker information to transform static text into dynamic portrait token with completion rhythm and amplitude information. To achieve coherence between segments, we design a Past Guided Generation Module (PGG) to maintain the consistency of customized listener attributes through the motion prior, and utilize a diffusion-based structure conditioned on the portrait token and the motion prior to realize the controllable generation. To train and evaluate our model, we have constructed two text-annotated listening head datasets based on ViCo and RealTalk, which provide text-video paired labels. Extensive experiments have verified the effectiveness of our model.
1 Introduction
Listening Head Generation (LHG) is to generate listener motions (face expressions, head movements, etc.) that respond synchronously to the speaker during face-to-face communication. Good communication is a two-way process, and the roles of speakers and listeners are equally important [4]. Whether the listener responds promptly will affect whether the speaker can actively express the follow-up content [17, 19].
Unlike speaker head generation (SHG) [7, 16, 31, 37], listeners can only convey information through non-verbal feedback (nodding, frowning, etc.), and their responses are affected by both sides of the interaction (speaker’s motion, tone, and listener’s own personality and emotion state). Currently, many applications, such as digital avatar generation [8, 20] and human-computer interaction [42], use the listener agent to simulate more realistic conversations, which has promoted many studies in LHG.
Figure 1: The process of text-guided listener generation in our CustomListener. The text prior provides the basic portrait style of the listener, which is input into CustomListener and combined with the speaker’s information to obtain the listener’s motions.
Some early works manually integrate motions based on rules [34, 1, 5], or rely on 2D facial keypoints [11, 26] to manipulate the head, but they are limited in the variety of motions and the detailed control ability. Recent works [43, 15, 21, 33, 25] have introduced 3D face coefficients [9], combined with the post-rendering module to achieve more precise control. Specifically, RLHG [43] uses a sequence-sequence structure as the baseline method for decoding 3DMM coefficients. PCH [15] further improves the render module to inpaint image artifacts. L2L [25] quantizes listener motions into the discrete one-dimensional codebook by VQ-VAE [35], and ELP [33] expands the codebook to a composition of several discrete codewords to achieve more fine-grained control. MFR-Net [21] designs a feature aggregation module to achieve better identity preservation and motion diversity.
Although above methods can combine speaker-listener information to generate diverse and fine-grained listener motions, they have restrictions in the explicit control of the listener agent’s response. Specifically, L2L [25] focuses on the motion diversity but cannot control the motion under specific conditions. RLHG [43], PCH [15], MFR-Net [21], and ELP [33] achieve controllable listener motion conditioned on the input attitude. However, the attitude is limited to a few fixed labels (eg. positive, neutral, negative), posing limitations on the generation of realistic listeners: First, simple labels are not enough to give accurate responses. For example, for negative attitudes, the two emotions of doubt and anger correspond to different facial expressions. More importantly, a realistic listener agent should have its own identity (parent, friend, teacher, etc.), personality (lively, calm, etc.) and behavioral habits (e.g. frowning when thinking), which users can customize and set in advance. However, simple labels cannot achieve the control under these conditions.
Considering above limitations of simple labels, in this paper, we propose a user-friendly framework called CustomListener to enable the text-guided generation, as shown in Figure 1. Users can pre-customize detailed attributes of the listener agent. Then using a large language model, we combine speech content and user-customized attributes to obtain the text prior of the listener’s basic portrait. In CustomListener, we seamlessly incorporate speaker information while being guided by the text prior, so as to generate realistic listener responses that are controllable and interactive.
For the guidance of text prior, due to the interactive nature of communication, listener motions cannot simply be regarded as text-conditioned motion generation, but faces the following challenges: 1) Speaker-listener coordination: the text provides the basic portrait style of the listener’s response. While to precisely express the listener’s emotional empathy towards the speaker, the response not only needs to complete the text-specified motion, but also needs to adjust its completion timing and rhythm, and fluctuate with the speaker’s semantics, intonation and movement amplitude. 2) Listener motion coherence: When motions conditioned on different text priors are combined into a long video, different segments should be coherent. Due to the customization, apart from ensuring smooth motion switching, we also need to guarantee that the user-customized listener’s behavioral habits reflected in past segments are maintained in the current segment. Maintaining the style of the past while still showing the motion of the current text can be challenging.
To solve above challenges, in our CustomListener, we design the Static to Dynamic Portrait module (SDP) and the Past-guided Motion Generation module (PGG) to achieve the coordination and the coherence respectively. In SDP, we transform the static portrait provided by the text prior to be dynamic one, which interacts with audio to obtain the time-dependent information about the completion time and rhythm of the gradual motion changes, and refine the listener’s motions to make the speaker-listener fluctuations relevant. In PGG, we generate a motion prior based on the similarity of dynamic portrait tokens between adjacent segments, which provides reference information for features that need to be maintained in the current compared to the past. Then, the motion prior and the dynamic portrait token are input as conditions into the diffusion-based structure [14] with fixed noise at the segment connection, finally generating listener coefficients that enable coordination and coherence in interactions.
In summary, this paper has the following contributions:
  • We argue that simple labels are not sufficient to achieve freely controllable listener generation, and thus propose the CustomListener framework that can freely control listener head motions guided by the text prior.
  • We design a SDP module to achieve the speaker-listener coordination, so that the completion timing and rhythm of text-specified motions correspond to the speaker, and fluctuate with the speaker’s semantics, tone, and movements.
  • For long-term generation, we design a PGG module to meet the listener motion coherence between different text-prior-conditioned segments, so that the consistency of the listener’s customized behavioral habits between segments can be maintained and the motion switching is smooth.
  • Extensive experiments on ViCo [43] and Realtalk [12] datasets confirm that our method can realize the state-of-the-art performance, and achieve the controllable and interactive realistic listener motion generation.
2 Related work
Figure 2: Overall framework of CustomListener. Given the text prior providing the listener’s static portrait style, SDP-Module transforms the static portrait into a dynamic one, which includes the time-dependent information and the fluctuation of motions. Then in PGG-Module, two adjacent portrait tokens are input to Past-guided Module to generate the motion prior, which is helpful to maintain the motion coherence. The portrait token and the motion prior are utilized as conditions of the diffusion-based structure to realize the controllable generation.
Listening head generation (LHG):
Some early works [34, 1, 5] use rule-based methods to manually incorporate interactive motions, which have limitations in motion diversity and generalization on the wild data. Later, data-driven methods [11, 26] generate 2D motions based on facial keypoints or movement frequencies, but 2D information cannot fully represent the details of facial movements.
Recently, motion generation based on 3D face coefficients, coupled with the post-processing renderer, has gradually become mainstream due to its stronger representation ability of facial movements. RLHG [43] proposes a representative ViCo dataset and constructs a baseline coefficient decoding method using a sequence-sequence structure. After that, PCH [15] assembles an enhanced renderer to make the generated results visually better and more stable by inpainting the image boundary and background. Learning2Listen (L2L) [25] proposes to use a sequence-encoding VQ-VAE [35] to learn the discrete codebook of listener motion. ELP [33] further utilizes a composition of multiple discrete motion-codewords to represent facial motion in a more fine-grained manner.
MFR-Net [21] achieves the preservation of listener’s identity and the diversity of generation based on the self-designed feature aggregation module.
Regarding the listener’s motion status, L2L [25] focuses on the diversity but cannot explicitly control specific actions. RLHG [43], PCH [15], MFR-Net [21] achieve controllable generation conditioned on the input emotion, but the emotion is limited to a few labels and cannot specifically control the expression and posture of each emotion. Although ELP [33] represents the latent space under different emotions in a fine-grained manner through the codebook, what the user can control is still the input emotion rather than the fine-grained motions. In this way, the listener’s response to each emotion may depend on the data distribution of the training set.
Conditional motion generation:
The motion generation task derives applications under a variety of conditions. For example, body motion conditioned on text descriptions [18, 27, 40, 28, 13], motion trajectory conditioned on scene image [3], and dance motion conditioned on music [38]. Recently, large language models promote the development of free-form text conditioned generation based on auto-encoder [28], hierarchical model [13], and diffusion models [18]. We also generate motions conditioned on free-form text. However, for our responsive interaction, in addition to focusing on the matching of motion and text like above methods, we also need to make the motion and the speaker’s action, speech, tone and rhythm harmonious to express the listener’s empathy. In our paper, we adopt the SDP module to achieve this harmony. For long-term generation, to ensure the coherent fluency between clips under different text commands, some methods have adopted transition sampling [39], specifying noise [41], etc.. However, the coherence for the custom listener, which has some motion habits, is more complex. It is not just a simple motion switch, but is related to the past habits and styles, and our PGG module solve this challenge.
3 Method
3.1 Overview
In our CustomListener framework, under the guidance of the fine-grained text prior, we can synthesize natural and controllable listener motions, which can seamlessly integrate the speaker’s information (motion, speech content, audio) to give interactive non-verbal responses. Let $L_{m}\in\mathbb{R}^{T\times 70}$ denote the listener motions to be generated, and $T$ is the number of frames, 70 denotes 64-dim expression and 6-dim pose. $R$ denotes the text prior. Given speaker’s motions $S_{m}\in\mathbb{R}^{T\times 70}$, audio $S_{a}$ and speech content $S_{t}$, our framework can be formulated as:
$L_{m}=Custom(S_{m},S_{a},S_{t};R)$ (1)
The overview of our proposed framework is shown in Figure 2. Before generation, for each video segment, we use GPT to generate a text prior based on user-customized listener attributes and the speaker’s speech content, which provides the listener’s static portrait reference. Then, we extract speaker’s audio features from the Audio Encoder and speaker’s motion coefficients from a pre-trained face reconstruction model [9]. Following this, we generate realistic listener motions through a two-stage process: (1) Static to Dynamic Portrait generation (SDP-Module): it can transform static portrait tokens into the dynamic through audio-text responsive interaction and speaker-motion based refinement. The dynamic portrait token includes the time-dependent information about the motion completion and rhythm, and makes the speaker-listener motion fluctuations correlated; (2) Past-guided Listener Motion Generation (PGG-Moudle): to produce coherent motions and maintaining the listener’s previous behavioral habits in long-term generation, we use a past guided module to produce the motion prior, combined with the diffusion denoising process with a fixed initial noise to generate the listener coefficients. Finally, listener videos are obtained through the post-rendering process.
3.2 Static to Dynamic Portrait generation
In this section, we first introduce the generation of the static portrait-token, and subsequently, we present SDP-Module, which can convert static tokens into dynamic one by two stages: (1) Audio-text Responsive Interaction: static portrait tokens are converted into time-dependent portrait tokens, which reflects the completion time and rhythm of motions according to the semantics and intonation in the audio, ensuring the gradual change of the motion; (2) Speaker Motion-based Refinement: it refines the amplitude of the listener’s motions according to the fluctuations of speaker’s motions, producing the final dynamic portrait token combined with the time-dependent portrait token.
Static Portrait-token Generation
To establish a better connection between the user-customized listener attributes (identity, personality, relationship, attitude, etc.) and the listener’s motions, we leverage the powerful GPT for semantic understanding of both user-customized contents and speaker’s speech contents, and generate the text prior that accurately describes the listener’s expressions and poses. Following this, we use RoBERTa [22], a robust and efficient model with high natural language understanding capability, to encode the text prior into text embeddings. To obtain portrait-related tokens, we further employ a mapping net to convert text embeddings into portrait tokens. Since the text prior describes the listener’s static portrait including expressions and head movements in the entire video clip, we consider the above portrait tokens generated from the text prior to be static portrait tokens.
Audio-text Responsive Interaction
Since static portrait tokens from text prior provide a comprehensive representation of listener head in a video segment, it lacks the time-dependent information when the motion is completed. Thus, the model would be limited to generating constant motion and cannot effectively convey gradual variation in movement without a specific design. To address this critical issue, we make use of the semantic knowledge in speaker’s utterances, which serves as a guidance for motion shifts. Specifically, we choose the speaker audio to provide semantics instead of speech contents. The underlying intuition is that the former contains not only semantic information but also the temporal information. This entire process can be splited into two parts, part A and part B.
Figure 3: Illustration of Audio-text Responsive Interaction. We first generate a weight matrix by responsive interaction between audio features and static portrait-token (SP Token), and then generate time-dependent portrait tokens (TP Token) guided by this weight matrix.
Part A In this part, we generate a weight matrix through responsive interaction between speaker audio features and the static portrait tokens obtained from text prior. It provides the correlation degree between the speaker’s semantics at each timestamp and the motion description in the text prior. For example, as shown in Figure 3, if the phrase “love you” in the speaker’s audio at time stamp $k$ exhibits the strongest correlation with the phrase “raised lip corner” in the description texts, the corresponding weight value will reach its peak (indicated by a pink circle dot in the weight map). This weight matrix $M(A,H)$ can be formulated as Equation 2.
Part B After obtaining a weight matrix $M(A,H)$, we can generate time-dependent portrait tokens based on static portrait tokens with the guidance of $M(A,H)$. Specifically, as illustrated in Figure 3, the time-dependent portrait tokens (TP Token) at time stamp $k$ is generated with the highest weighting assigned to the static portrait tokens (SP Token) at $l=j$, which corresponds to the phrase “raised lip corner”. This part can be formulated as Equation 3.
$M(A,H)=Sm(\frac{1}{\sqrt{\alpha}}(Linear(A)\times Linear(H)^{T})),$ (2)
$E_{t}=\sum_{i=0}^{L-1}M(A,H)_{i,t}\cdot H_{i,\operatorname{j}=0:C-1},t\in\left\{0,...,T-1\right\}$ (3)
where $Sm$ denotes Softmax, $A\in\mathbb{R}^{T\times C}$ is the audio feature, $H\in\mathbb{R}^{L\times C}$ is the static portrait token, $M\in\mathbb{R}^{T\times L}$ is the weight map, $\alpha$ is a scaling factor, $E_{t}\in\mathbb{R}^{1\times C}$ is the time-dependent portrait token at time stamp $t$. $L$ is the number of word segments in the text, and $C$ is the feature dimension.
With the help of two parts above, we can obtain portrait tokens that contain time-dependent information about the rhythm of motion completion. This enables the model to learn the progressive changes in facial/head motions within a single video segment, as opposed to the constant motions.
Speaker Motion-based Refinement
In human-to-human conversations, we believe that the amplitude of the speaker’s motion fluctuation also has an impact on the changes in the listener’s motion response. Therefore, we use the cumulative value of the speaker’s motion changes in the past window period to refine the listener’s motion amplitude. This process can be well-designed formulated as:
$\overline{E_{t}}=\sum_{i=t-1-w}^{t-1}\exp\left|\frac{\partial S_{i}}{\partial i}\right|,$ (4)
$\widetilde{E_{t}}=E_{t}\oplus\overline{E_{t}},$ (5)
where $\oplus$ means concatenation, $E_{t},\overline{E_{t}},\widetilde{E_{t}}$ is time-dependent portrait token, speaker motion-weighted portrait token and dynamic portrait token at frame $t$, respectively. $S_{i}$ denote speaker motion at frame $i$, $w$ is the time-window length. For efficiency, we set $w$ to 5, which means we only take the past five-frame speaker motions into account. For each time stamp $t$, the listener’s portrait token $E_{t}$ is affected by the cumulative gradients of several previous speaker motions. This process can be illustrated in the lower part of the SDP-Module in Figure 2. Notably, when generating portrait token at frame $t$, we only focus on the speaker motions up to frame $t-1$. Due to the listener and the speaker communicating concurrently, the listener’s state at frame $i$ should not influenced by the speaker’s motion at the same frame $i$.
3.3 Past-guided Listenr Motion Generation
The PGG-Module is to ensure that motions between segments are coherent, and the customized listener attributes reflected in the previous segment can be maintained in the current segment. It contains Past-guided Module and diffusion-based motion generation module, which can generate motion prior conditioned on past motions and then synthesize natural listener motions. Following this, we employ a pretrained render model [30] to obtain final listener videos.
Past-guided Module
To produce long video with arbitrary length, we opt for a segment-by-segment generation approach during inference. However, we have observed that there are unsmooth pose transitions between segments. This issue can be attributed to the neglect of correlations between adjacent segments. To address this, we take past motions into consideration. Since our final goal is to synthesize a customized listener, in addition to ensuring smooth transition, we also need to take the listener’s past behavioral habits into account (e.g. frowning when thinking). Hence, we sought assistance from dynamic portrait tokens of two adjacent segments to guide the generation of motion prior.
Specifically, the weight we assign to past motions at different times is determined by the similarity between portrait tokens in the current segment and those in the last segment. Formally, we calculate the motion weight and obtain the motion tokens by:
$Sim_{t:t+L}=E_{t:t+L}\times E_{t-L:t}^{T},$ (6)
$G_{t:t+L}=Sim_{t:t+L}\times m_{t-L:t},$ (7)
where $\times$ denotes matrix multiplication, $L$ is the segment length, $t$ is the time stamp, $G_{t:t+L}$ is the generated motion prior, $Sim_{t:t+L}$ denotes the similarity between portrait tokens in current segment $E_{t:t+L}$ and those in last segment $E_{t-L:t}$, which is the weight we assign to past motions $m_{t-L:t}$. The higher the similarity, the greater the weight we assign. This approach can not only encourage continuous transitions between two adjacent segments, but also ensure the maintenance of customized listener’s behavioral habits between two adjacent sequences.
Diffusion Model
To achieve diverse and natural head motion generation, our generation module is built on denoising diffusion probabilistic models (DDPM) [14]. And thus we can fix the initial noise of current segment to be the end of the previous segment to ensure smooth motion switching.
During forward diffusion process $q$, we progressively inject noise into the listener motion $x_{0}$ until it becomes a Gaussian noise $x_{t}$, which can be defined as a Markov chain:
$q(x_{1:T}\mid x_{0})=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$ (8)
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (9)
where $x_{1},...,x_{t}$ are noised listener motions, $\beta_{1},...,\beta_{t}$ are the variance schedule. In the reverse process $p_{\theta}$, we produce listener motion $x_{0}$ by removing noise from $x_{t}$ step-by-step:
$p_{\theta}(x_{t-1}\mid x_{t},c)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},c,t),{\textstyle\sum_{\theta}}(x_{t},c,t)),$ (10)
where $c$ is a conditioning variable. We perform denoising process with a transformer decoder [36] conditioned on the motion prior and dynamic portrait token.
Render
We utilize PIRenderer [30] for rendering, which can produce natural listener videos by incorporating a single listener reference image as well as the 3DMM coefficients.
4 Experiments
[…]
exp angle trans exp pose exp pose exp pose exp pose exp pose
${}^{\ast}$ [43] $\mathcal{D}_{test}$ 15.03 7.90 6.55 0.113 0.160 0.108 0.160 11.65 0.90 2.90 0.09 0.69 0.10
$\mathcal{D}_{ood}$ 22.81 8.58 8.90 0.124 0.101 0.104 0.100 9.32 0.82 7.20 0.11 1.01 0.12
${}^{\ast}$ [15] $\mathcal{D}_{test}$ 19.02 13.26 7.74 0.124 0.163 0.109 0.198 11.59 0.96 3.10 0.13 0.31 0.12
$\mathcal{D}_{ood}$ 18.63 17.96 8.80 0.121 0.115 0.107 0.140 8.67 0.87 7.01 0.12 0.51 0.16
${}^{\ast}$ [25] $\mathcal{D}_{test}$ 14.89 7.35 6.34 0.102 0.153 0.101 0.113 8.64 0.87 2.62 0.09 0.57 0.20
$\mathcal{D}_{ood}$ 15.64 8.23 7.89 0.095 0.098 0.100 0.101 6.78 0.74 6.89 0.10 1.23 0.28
${}^{\dagger}$ [21] $\mathcal{D}_{test}$ 13.37 6.82 6.02 - - - - - - - - - -
$\mathcal{D}_{ood}$ 14.70 8.12 6.37 - - - - - - - - - -
Ours $\mathcal{D}_{test}$ 11.54 6.12 5.90 0.072 0.103 0.081 0.097 3.12 0.06 2.40 0.07 1.55 0.32
$\mathcal{D}_{ood}$ 12.67 7.49 6.01 0.081 0.073 0.084 0.075 3.59 0.07 5.94 0.05 1.25 0.30
[…]
4.1 Datasets
Data Annotation
Since the existing listener head datasets do not contain the full data required for training our model, we conduct additional data annotations on two datasets:(1) ViCo [43], a popular dataset in listener head generation; (2) RealTalk, proposed by [12], which serves as a database for retrieving listener videos in [12]. Firstly, to learn the relationships between the text prior and listener expressions, we follow the text annotation pipeline in TalkCLIP [24], and conduct text annotation on ViCo and RealTalk. Specifically, we employ [6] to acquire emotion labels, activated AUs as well as their intensities. Additionally, to learn realistic head movements such as nodding to convey agreement and shaking the head to signify disagreement, we use Hopenet [32] to detect head motions. Finally, we generate diverse text prior description for each video segment according to the format: [A person $<$EMOTION$>$ and listens with $<$AU$>$ (and $<$HEAD MOTION$>$)], where $<$EMOTION$>$ denotes the emotional labels, $<$AU$>$ denotes fine-grained expression-related texts based on AU activation intensities and Facial Action Coding System [10], and $<$HEAD MOTION$>$ indicates head nodding/shaking labels. We use “and” to link when there is more than one activated AU.
Data Construction
For ViCo, which contains three parts: train set $\mathcal{D}_{train}$, test set $\mathcal{D}_{test}$, out-of-domain set $\mathcal{D}_{ood}$ (all identities in $\mathcal{D}_{ood}$ have not appeared in $\mathcal{D}_{train}$), we obtain the transcripts of the speaker’s audios by a speech recognition model [29], and get an extended version of ViCo. For RealTalk, to align the dimension of 3DMM coefficients with ViCo, we re-extract facial 3DMM coefficients using [9] and then acquire two subsets: $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$. These two extended datasets are important for future research on text-guided listener head generation task. Notably, we did not use L2L [25] datasets as we cannot obtain the raw videos of L2L [25] and thus cannot perform text annotations on it.
Figure 4: Visual results produced by CustomListener. All listener videos are generated conditioned on different pre-set text priors, the same speaker video (the 1st row) and the same reference listener image.
4.2 Experimental Settings
Loss Fuction
During training, followed by [14], we perform denoising on noised listener motions $x_{t}$ conditioned on motion prior and dynamic portrait token step-by-step, and our optimization goal is to predict noise term $\epsilon$ at each step. The objective function can be defined as a conditional diffusion loss as follows:
$\vspace{-0.1cm}L_{d}=E_{x_{0},t,\epsilon_{t}\sim\mathcal{N}(0,1)}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(x_{t},c,t)\right\|^{2}\right],\vspace{-0.1cm}$ (11)
where $\epsilon_{t}\sim\mathcal{N}(0,1)$ is the noise at step $t$, $c$ denotes conditions, $x_{0}$ is the original listener motions.
Implementation Details
The resolution of facial videos is 256x256, and the FPS is 30. For training, we clip the original videos to several 60-frame video segments. We extract 3DMM coefficients from speaker videos and listener videos by [9], which serve as speaker motions and listener motions, respectively. Following MFR-Net [21], we extract 45-dim acoustic features from audios, containing 14-dim MFCC, 28-dim MFCC-Delta, energy, Zero Crossing Rate and loudness. We utilize AdamW optimizer [23] to train our model with a learning rate of $1e^{-4}$.
Method SSIM $\uparrow$ CPBD $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
${}^{\ast}$ [43] 0.56 0.12 17.39 27.70
${}^{\ast}$ [15] 0.58 0.15 18.48 21.29
${}^{\dagger}$ [21] 0.59 0.18 17.82 20.08
Ours 0.60 0.19 17.98 20.06
Table 2: Quantitative results with state-of-the-art methods on image quality. The best results are highlighted in bold.
4.3 Quantitative Results
Metrics
It’s non-trivial to measure the realism of generated listener motions. We choose the following metrics based on that a good custom listener should have four characteristics: (1) natural-looking and diverse facial/head motions; (2) highly synchronous with the speaker motions; (3) highly coherent motions in a long video; (4) high image quality after rendering:
$\bullet$ FD: $L1$ distance between the predicted 70-dim listener motions and the ground-truth listener motions.
$\bullet$ Variation for Diversity (V-D) : Variance across the time series sequence of expression coefficients or pose coefficients.
$\bullet$ Residual Time Lagged Cross Correlation (RTLCC): $L1$ distance between generated TLCC [2] and TLCC [2] of ground truth. RTLCC indicates the correlation between listener motions and lagged speaker motions.
$\bullet$ Residual Windowed Time Lagged Cross Correlation (RWTLCC): RTLCC in a fixed window. We set the window length to 4 seconds following ELP [33].
[…]
$\bullet$ SND: Sequence Naturalness Distance proposed by [41], showing the distribution difference between generated and ground-truth motions.
$\bullet$ SSIM, CPBD, PSNR and FID: Common metrics for image quality evaluation.
Figure 5: Qualitative comparisons with PCH [15], RLHG [43] based on the same speaker and the same listener reference image.
Comparisons with Existing Methods
We retrain PCH [15], RLHG [43] and L2L [25] on VICO dataset. It should be noted that, due to the unavailability of source codes from ELP [33] and MFR-Net [21], for MFR-Net [21], we utilize the data from the original paper. For ELP [33], the 3DMM coefficient extraction model used is different from ours, resulting in distinct dimensions of coefficients (e.g., $\beta\in\mathbb{R}^{100T}$ for ELP [33], $\beta\in\mathbb{R}^{64T}$ for ours), thus it is not reasonable to directly compare with evaluation data in ELP [33]. Therefore, we only provide its visual comparisons in the Appendix C. Our quantitative comparisons consist of four parts: 1) precision and diversity of listener motions: as presented in Table 1, for the realism of listener motion generation, our proposed CustomListener achieves lowest FD on both $\mathcal{D}_{test}$ and $\mathcal{D}_{ood}$ subsets in ViCo. Meanwhile, the V-D of our model is comparable to other methods, which indicates that our model can maintain certain diversity when generating controllable listener motions. This may benefit from the probabilistic nature of the diffusion model. 2) speaker-listener synchronization: compared to all existing methods, CustomListener can decrease the average RTLCC and the RWTLCC,
[…]
Method FD $\downarrow$ RTLCC $\downarrow$ RWTLCC $\downarrow$ FID ${}_{\triangle fm}$$\downarrow$ SND $\downarrow$
(x100) (x100)
w/o ATRI 20.67 0.206 0.203 5.01 4.53
w/o SMW 19.05 0.201 0.196 5.07 4.58
w/o SDP 21.97 0.213 0.210 5.90 4.60
w/o PGM 20.53 0.197 0.201 6.47 4.65
Ours 18.48 0.165 0.169 3.42 4.23
Table 3: Ablation study. Each cell represents the average results of 70-dim motions (64-dim exp + 6-dim pose) on $\mathcal{D}_{test}$ and $\mathcal{D}_{ood}$.
Figure 6: Ablation study of PGM to verify that PGM can maintain motion coherency. Each group includes adjacent four frames, and the frames enclosed by the red box are the two frames at the transition point between two video clips.
4.4 Qualitative Results
Visual Results
Since previous methods cannot generate listener head based on fine-grained texts, we first present visual results generated by CustomListener. As shown in Figure 4, given different texts, covering diverse listener’s personalities, identities and attitudes toward the speaker’s utterances, our framework can produce fine-grained non-verbal listener motions closely aligned with pre-customized texts. For example, if the listener is an impatient person who is not interested in the speaker’s words(the 3rd row in Figure 4), he may appear indifferent and show impatience. Additionally, the produced listener’s responses can fluctuate with the speaker’s semantics and movement, showing great effectiveness of our proposed dynamic portrait tokens.
Comparisons with Existing Methods
In Figure 5, we compare our results with RLHG [43], PCH [15]. It can be seen that our method is capable of generating more realistic listener heads than other baselines. In the left four columns, both RLHG [43] and PCH [15] suffer from unnatural motions (e.g., disorderly opened mouth) and visible facial artifacts. In the right four columns, RLHG [43] still exhibits incorrect motions and PCH [15] can not display negative attitude accurately. While our methods can produce natural listener head that highly aligns with the ground truth conditioned on the given texts. For more detailed image/video format comparisons, please refer to our Appendix C.
Figure 7: Ablation study of PGM. For different segments, the habit of frowning is maintained using PGM.
4.5 Ablation Study
SDP module
In Section 3.2, we design a SDP module to generate dynamic portrait tokens from static one for the purpose of realizing progressive motion shifts that correspond well to the speaker. To validate its effectiveness, we conduct ablation study on Audio-text Responsive Interaction (ATRI) and Speaker Motion-based Weighting (SMW), respectively. The results are shown in Table 3. In addition to the improvement in FD, the improvement in the RTLCC and RWTLCC is more pronounced, which demonstrates that the SDP module is crucial for enhancing speaker-listener synchronization, and each of its components matters.
Past-guided module
[…]
5 Conclusion
In this paper, we present a user-friendly framework called CustomListener, which can realize freely-controllable listener head generation conditioned on text priors. To achieve the speaker-listener coordination, we propose a SDP-module for dynamic portrait-token generation, which is capable of representing progressive motion changes. To ensure coherent listener responses in a long video, we design a PGG-module for smooth transition between adjacent video segments as well as maintaining customized behavioral habits. Comprehensive experiments demonstrate the superiority of our whole framework. In future work, it may be interesting to generalize our framework to listening body generation, yielding a complete and realistic listener agent.